### v1

@@ Results
Similar patterns exist in the public health domain, where conspiracy beliefs, including those about international elites, tend to align with anti-vaccination attitudes (Hornsey, Harris and Fielding 2018). While these were fringe positions before the COVID-19 period, they became more mainstream during the pandemic. On social media, anti-vaccination and anti-internationalist conspiracy theories became rampant (Darius and Urquhart 2021; Erokhin, Yosipof and Komendantova 2022). Offline, anti-vaccine individuals in the U.K. often simultaneously held pseudo-scientific, populist, and conspiracy beliefs (Holford et al. 2023), and a sizable segment of the U.S. population negatively viewed the World Health Organization (WHO) during the pandemic (Bayram and Shields 2021).
Did the COVID-19 pandemic elevate existing populist sentiments against climate science and policies? Based on prior research showing that cognitively coherent beliefs reinforce one another (Taber and Lodge 2006), we expect overlapping antagonistic populist beliefs toward international elites and scientists in the realms of public health and climate change to have served as a pathway for spillover between the two global crises, leading to greater antagonism toward international organizations in climate discussions during the pandemic.
@@
Table 3: Examples of climate tweets with high probability of toxicity (above 0.75) that mention international organizations before and after the onset of COVID-19 pandemic. Tweet texts have been edited for typos and formatting.
To explore this possibility, we examined the temporal dynamics, before and after the pandemic’s onset, of toxicity differences between climate tweets that referenced international organizations and those that do not. Because anti-internationalist sentiments lead to greater incivility in climate discussions, if the anti-internationalist sentiments activated by COVID-19 spilled into the climate domain, there should be greater toxicity in climate change discussions about international organizations compared to the baseline, even in tweets not directly concerning the pandemic. As the onset of the pandemic was exogenous to climate change in the short term, observed temporal differences can be attributed to the COVID-19 shock.
@@
Looking at second differences in a linear probability model of tweet toxicity (Mize 2019), we find a significant increase in toxicity toward international organizations in climate discussions during the COVID-19 period starting Q1 2020, when compared to the baseline period of Q4 2019 (Figure 3). The difference in toxicity probability before the pandemic, on the other hand, was relatively stable. The highest spike occurred in Q2 2020, immediately following the WHO declaring COVID-19 to be a global pandemic, and not until Q3 2021 did the difference fall to pre-pandemic levels. These findings provide clear evidence that the nature of climate discussions about international organizations changed with the pandemic.
@@ Discussion
While the COVID-19 pandemic and its public salience has subsided, legacies of pandemic era politics continue today. Science being contested as part of the political process, which was apparent with the political vitriol surrounding Anthony Fauci, continues with politically-motivated appointments to science-related posts following government transitions (Tanne 2024). Trust in scientists remains relatively low at levels similar to the pandemic period (Pew Research Center 2024), and recent research shows worrying signs of behavioral spillover between COVID-19 and post-pandemic vaccination attitudes (Lunz Trujillo et al. 2024). With climate change remaining ever as important, it is imperative to better understand how spillover among concurrent crises might lead to system-wide entrenchment of affectively polarized attitudes.
@@
In this study, we showed that affective polarization surrounding COVID-19 spilled over to how the public engages with climate change and climate science on social media. References to COVID-19, especially those surrounding Anthony Fauci, were strongly associated with incivility and contentiousness in climate change discussions. With some exceptions, COVID-19 content both invited and were used for antagonistic engagement with the climate issue. These relationships were responsive to landmark pandemic era events, as incivility and contentiousness increased when the role of science in policymaking was made more salient by COVID-19 policies. We further found evidence that this spillover occurred in part through existing sources of affective polarization. Populist sentiments against international organizations and associated elites provided a link between climate policy opposition and COVID-19 vaccine hesitancy, which manifested in higher incivility of climate discussions even when COVID-19 was not referenced.
@@
By demonstrating the existence of affective polarization spillover and how it activated along existing political cleavages, our study adds to the literature on how COVID-19 impacted the public’s engagement with the climate issue, which has thus far focused on how concurrent crises can lead to shifts in their relative issue saliency (Smirnov and Hsieh 2022; Sisco et al. 2023) and the potential for policymaking synergies between them (Bergquist et al. 2023). More broadly beyond COVID-19 and climate change specifically, our findings show that conceptual synergies can lead to harmful antagonistic behavior spilling over across different domains, especially for public policy issues that are strongly informed by science, such as public health and environmental governance, highlighting the dangers of entrenched affective polarization on how the public engages with science.
Data and Methods
@@
We used several data sets of online social media posts published between February 1, 2019 and August 26, 2021, spanning a total of 134 weeks. We subsetted all posts to English ones using the language estimate returned by Perspective API. We collected all tweet data using the full archive search endpoint from Twitter’s v2 API suite during October 2021–June 2022. Additional Twitter data collection details are in Section S1.
General climate tweets
38.6 million tweets from 5.6 million unique users published in the data collection window containing at least one climate change related keyword or hashtag.
General climate conversations
Full reply tree of 311,000 conversations (stratified random sample of approximately 2300 per week) from the General climate tweets data set that have at least one reply and two unique participants.
Climate scientific tweets
265,000 tweets from 81,000 unique users containing a reference (hyperlink) to a scientific publication containing climate science related keywords in the publication title. This data set was curated using Altmetric’s API made available through its Researcher Data Access Program.
Climate scientific conversations
Full reply tree of all 42,000 conversations from the Climate scientific tweets data set that have at least one reply and two unique participants.
Reddit climate change comments
2.1 million Reddit comments mentioning the keyword “climate change”. This data set was provided by SocialGrep (2022). Preprocessing details are in Section S2.
@@ 4.2.1 Topic Classification
Based on the presence or absence of specific keywords, all posts were assigned binary labels for each of the following categories: 1) For COVID-19, we used keywords and hashtags generally used to refer to the virus or pandemic and its interventions (e.g., vaccines, social distancing); 2) for Anthony Fauci, we only used his surname, as it is distinct enough to capture relevant references to him and discriminate against false positives; 3) for international organizations, we used a list of keywords that include international entities (i.e., international organizations and public figures). The full keyword lists are in Section S3.
@@ 4.2.3 Disagreement
To understand the level of disagreement among participants, we fine-tuned a 3-class classifier based on the DistilBERT language representation model (Sanh et al. 2019) using the DEBAGREEMENT comment-reply agreement and disagreement data set (Pougué-Biyong et al. 2021), a set of 43 thousand comment-reply pairs from political and climate subreddits, each labeled as either disagreement, neutral, or agreement. Because we are only concerned with the presence of disagreement, constructing a binary classifier based on the three class probabilities results in $f_{1}$ scores of 0.806 and 0.716 for disagreement and lack of disagreement (i.e., neutral or agreement) respectively. Performance details of our fine-tuned DistilBERT model are in Section S4.
@@
Linear Probability Model and Fractional Logistic Regression Model
We fitted all models in R (R Core Team 2024) with the fixest package (Bergé 2018). Confidence intervals were obtained using the marginaleffects package (Arel-Bundock, Greifer and Heiss 2024).
@@
We fitted separate linear probability models of COVID-19 period (March 20, 2020–August 26, 2021) climate post toxicity for all three of our data sets – general climate tweets, climate science tweets, and climate Reddit posts. In all three models, we included day fixed effects. We additionally included account fixed effects for the general climate tweets data set in a separate model. In all models, we clustered the standard errors at the day level. Fitting all models using fractional logistic regression (Papke and Wooldridge 2008) yielded substantively similar results, which we show in Section S6.
@@
We fitted separate linear probability models of COVID-19 period response disagreement for both of our climate Twitter data sets. In both models, we included day fixed effects, and clustered the standard errors at the day and conversation levels. Results from fractional logistic regression models are substantively similar, which we show in Section S6.
@@
We fitted a model for tweet toxicity using our general climate Twitter data set for our entire observation period (February 1, 2019–August 26, 2021). In the model, we interacted the international organization keyword term with every quarter in the data (Q1 2019–Q3 2021), then compared the coefficient of the international organization term from each quarter to the pre-pandemic baseline (Q4 2019). We computed standard errors both with and without account clustering. To ensure the increased incivility during the pandemic is not directly due to COVID-19 tweets, we removed them from this analysis, but our results are robust to whether these tweets are included, which we show in Section S8.
Cox Proportional Hazards Model
To understand the effect of mentions of COVID-19 and other related topics on the onset of toxicity in the course of climate conversations, we constructed an ensemble of Cox proportional hazard models in R (R Core Team 2024) using the survival package (Therneau 2024). Specifically, we modeled the first occurrence of a toxic reply in a Twitter thread, or consecutive chain of tweets and replies, from the COVID-19 period. The analysis is based on a resampling method. In each realization, one thread from each conversation tree is selected at random to ensure tweets closer to the root of the conversation are not over-represented in the analysis, as they appear in more threads than tweets closer to the leaves.
@@
Full details of how we implemented the Cox proportional hazards model are in Section S5. We show that our results are robust to alternative specifications of what constitutes an English conversation in Section S7.
@@
We used Twitter’s Academic API to collect our Twitter data. To collect the general climate tweets, we queried the API for tweets with hashtags and keywords from Table S3.1 for the February 1, 2019–August 26, 2021 period (missed data from February 4, 2021). We subsetted the data to only English tweets using Perspective API’s language estimate. We also excluded tweets containing “political climate” and “Percip:”. The latter is frequently used by bots that automatically tweet regional weather forecast. This yielded 38.6 million tweets.
To identify all tweets that reference (through hyperlinks) climate science publications, we first used Altmetric’s Explorer API to filter scientific publications by whether their titles contained one of our keywords (Table S3.1), then used Altmetric’s Details Page API to obtain IDs of all tweets that referenced these publications. We used Twitter’s Academic API to collect these tweets, which we subset to only English ones. This yielded 265,000 tweets.
@@
Table S1.2 shows the distribution of the number of users that were active before and after the COVID-19 onset.
@@
Only during the pandemic 2.2M 38K
@@
For our Reddit analysis, we used the Reddit Climate Change Dataset provided by SocialGrep (2022). This data set consists of approximately 4.6 million Reddit comments posted before September 2022 that contain “climate change”. We removed comments posted outside our time window and non-English comments. We also removed comments containing phrases indicating that they were written by bots. Specifically, we filtered the comments on a set of common phrases used by bots on Reddit to perform automatic tasks, e.g. “I am a bot” or “This action was performed automatically” (Table S2.1), as well as excluding all comments from the subredditsummarybot because it mostly consist of automatic submissions. Our final Reddit data set contains 2.1 million comments.
Table S2.1: Phrases used for filtering automatic Reddit comments.
@@
Table S3.1: Keywords used for determining mentions of different topics in tweet texts
@@
We analyzed the general climate conversations and climate science conversations separately. For both sets of analysis, we used all English conversations March 20, 2020 onward from the respective data sets.
@@
  • Tweets with toxicity probability that are not able to be estimated by Perspective API [^3] Tweets in languages not supported by the Perspective API (i.e., those outside of Arabic, Chinese, Czech, Dutch, English, French, German, Hindi (including when written with Latin alphabet), Indonesian, Italian, Japanese, Korean, Polish, Portuguese, Russian, Spanish and Swedish), and tweets containing only non-textual material (e.g., images, videos, or links). Perspective API returned NA values for 3.4% of general climate tweets and 6.5% of climate science tweets. and all subsequent replies in its thread are removed, because we are unable to assess how they affect subsequent replies.
@@
As threads from the same conversation are not independent from one another due to sharing tweets higher up in the conversation (i.e., closer to the root tweet), we based our analysis on a resampling method. In each realization, we randomly sampled one thread from each conversation. This means that no two threads in the same realization come from the same conversation, and therefore do not shared any tweets. This resampling eliminated the dependence between chains from the same conversation, and also ensured that the proportional hazard assumption is not violated due to presence of extremely large conversation.
@@
Table S5.1: The proportion of bootstraps satisfying Cox proportional hazard model assumptions by not failing the Schoenfeld test with $p>0.05$.
@@
Fauci keywords 0.99 1.00
International root 0.98 1.00
International keywords 1.00 0.95
@@
Linear probability models often provide an adequate approximation of non-linear models, especially when the range of outcome probabilities are not extreme. Still, we fit additional fractional logistic regression models (Papke and Wooldridge 2008) of toxicity probability and disagreement probability with the same model specifications as what we used for our main analysis (Figure S6.1) to show that our results are robust to accounting for the bounded nature of probabilities.
Figure S6.1: Robustness check of main results using fractional logistic regression models. Panel A shows the average marginal effect estimates for toxicity probability, corresponding to results from Figure 1.A. Panel B shows the average marginal effect estimates for disagreement probability, corresponding to Figure 2.A.
@@
Figure S7.1: Robustness check of main Cox proportional hazards model results using full replies instead of only English ones, corresponding to Figure 1.B.
@@
When assessing the impact of COVID-19 on the relationship between anti-internationalist sentiment and toxicity in climate discussions, we examined only tweets that do not directly engage with COVID-19 to ensure our results are not driven purely by the toxicity of COVID-19 content. Still, we show that our results are robust to the inclusion of COVID-19 tweets (Figure S8.1).
Figure S8.1: Robustness check of toxicity probability of international organization tweets over time when including tweets with overlapping COVID-19 content, corresponding to Figure 3.

### v2

@@ 2.2.1 Disagreement and Climate Obstructionist Content
Table 3: Examples of uncivil climate tweets that mention international organizations before and after the onset of COVID-19 pandemic. Toxicity probabilities of these tweets exceed 0.75. Tweet texts have been edited for typos and formatting.
Similar patterns exist in the public health domain, where conspiracy beliefs, including those about international elites, tend to align with anti-vaccination attitudes (Hornsey, Harris and Fielding 2018). While these were fringe positions before the COVID-19 period, they became more mainstream during the pandemic. On social media, anti-vaccination and anti-internationalist conspiracy theories became rampant (Darius and Urquhart 2021; Erokhin, Yosipof and Komendantova 2022). Offline, anti-vaccine individuals in the U.K. often simultaneously held pseudo-scientific, populist, and conspiracy beliefs (Holford et al. 2023), and a sizable segment of the U.S. population viewed the World Health Organization (WHO) negatively during the pandemic (Bayram and Shields 2021).
Did the COVID-19 pandemic elevate existing populist sentiments against climate science and policies? Based on prior research showing that cognitively coherent beliefs reinforce one another (Taber and Lodge 2006) – which has been demonstrated in belief systems of conspiracy theories (Enders et al. 2021) – we expect overlapping antagonistic populist beliefs toward international elites and scientists in the realms of public health and climate change to have served as a pathway for spillover between the two global crises, leading to greater antagonism toward international organizations in climate discussions during the pandemic.
To explore this possibility, we examined the temporal dynamics, before and during the pandemic, of toxicity differences between climate tweets that referenced international organizations and those that do not. Because anti-internationalist sentiments lead to greater incivility in climate discussions, if the anti-internationalist sentiments activated by COVID-19 spilled into the climate domain, there should be greater toxicity in climate change discussions about international organizations compared to the baseline, even in tweets not directly concerning the pandemic. As the onset of the pandemic was exogenous to climate change in the short term, observed temporal differences can be attributed to the COVID-19 shock.
@@
Looking at second differences in a linear probability model of tweet toxicity  (Mize 2019), we find a significant increase in toxicity toward international organizations in climate discussions during the COVID-19 period starting Q1 2020, when compared to the baseline period of Q4 2019 (Fig. 3). The difference in toxicity probability before the pandemic, on the other hand, was relatively stable. The highest spike occurred in Q2 2020, immediately following the WHO declaring COVID-19 to be a global pandemic, and not until Q3 2021 did the difference fall to pre-pandemic levels. These findings provide clear evidence that the nature of climate discussions about international organizations changed with the pandemic.
@@ Discussion
While the COVID-19 pandemic and its public salience has subsided, legacies of pandemic era politics continue today. Science being contested as part of the political process, whether it be climate change or public health, was apparent with the political vitriol surrounding Anthony Fauci. It continues with politically-motivated appointments to science-related posts following government transitions (Tanne 2024). Trust in scientists remains relatively low at levels similar to the pandemic period (Pew Research Center 2024), and recent research shows worrying signs of behavioral spillover between COVID-19 and post-pandemic vaccination attitudes (Lunz Trujillo et al. 2024). With climate change remaining ever as important, it is imperative to better understand how spillover among concurrent crises might lead to system-wide entrenchment of affectively polarized attitudes.
@@
In this study, we showed that affective polarization surrounding COVID-19 spilled over to how the public engages with climate change and climate science on social media. References to COVID-19, especially those surrounding Anthony Fauci, were strongly associated with incivility and contentiousness in climate change discussions. With some exceptions, COVID-19 content both facilitated and was used for antagonistic engagement with the climate issue. On the whole, these relationships were responsive to landmark pandemic era events, as incivility and contentiousness increased when the role of science in policymaking was made more salient by COVID-19 policies. Still, while our observations generally held across all social media systems examined, there are differences between general climate discussions and discussions explicitly engaging with climate science. Whereas COVID-19 primarily spilled into general climate discussions as increased incivility, it entered science-focused discussions in the form of higher contentiousness but rarely devolved into personal attacks on scientists.
We further found that populist sentiments against international organizations and associated elites provided a link between climate policy opposition and COVID-19 vaccine hesitancy, which manifested in higher incivility of climate discussions even when COVID-19 was not referenced. This provides evidence that antagonism spillover occurred in part through existing sources of affective polarization, which supports our understanding that this spillover is part of a larger political sorting process that produced entrenched belief systems governing belief formation and political communication.
@@
By demonstrating the existence of affective polarization spillover and how it activated along existing political cleavages, our study adds to the literature on how COVID-19 impacted the public’s engagement with the climate issue, which has thus far focused on how concurrent crises can lead to shifts in their relative issue saliency (Smirnov and Hsieh 2022; Sisco et al. 2023) and the potential for policymaking synergies between them (Bergquist et al. 2023). More broadly beyond public health and climate change specifically, our findings suggest that conceptual synergies rooted in populist sentiments can lead to harmful antagonistic behaviors spilling over across different public policy domains that are strongly informed by science – such as environmental governance, foreign aid, and science funding – reiterating the dangers of entrenched group cleavages on how the public engages with issues at the science-policy nexus.
Materials and Methods
@@
We used five data sets, summarized in Table 4, of online social media posts published between February 1, 2019 and August 26, 2021, spanning a total of 134 weeks. We collected all tweet data using the full archive search endpoint from Twitter’s v2 API suite during October 2021–June 2022. To ensure consistency in our measures, most of which rely on natural language processing, we only used English tweets, i.e., those that were determined to contain English by both Twitter’s and Perspective’s language classification. Additional Twitter data collection details are in supplementary text.
Data Set Description Sample Size Users
General climate tweets Tweets published in the data collection window containing at least one climate change related keyword or hashtag 38.4M 5.5M
Climate science tweets Tweets containing a hyperlink to a scientific publication containing climate related keywords in its title. Data set curated using Altmetric’s API. 258K 79K
Reddit climate Comments containing the phrase “climate change”, labeled by Perspective to contain English. Data set provided by SocialGrep (2022). 2.1M
General climate conversations Full reply tree of 287,000 conversations (stratified random sample of approximately 2100 per week) from the General climate tweets data set that have at least one reply and two unique participants. 1.9M
Climate science conversations Full reply tree of all 38,000 conversations from the Climate science tweets data set that have at least one reply and two unique participants. 181K
Table 4: Summary of all data sets used.
@@ 4.2.1 Topic Classification
Based on the presence or absence of specific keywords, all posts were assigned binary labels for each of the following categories: 1) For COVID-19, we used keywords and hashtags generally used to refer to the virus or pandemic and its interventions (e.g., vaccines, social distancing); 2) for Anthony Fauci, we only used his surname, as it is distinct enough to capture relevant references to him and discriminate against false positives; 3) for international organizations, we used a list of keywords that include international entities (i.e., international organizations and public figures). The full keyword lists are in Table S3.1.
@@ 4.2.3 Disagreement
To understand the level of disagreement among participants, we fine-tuned a 3-class classifier based on the DistilBERT language representation model (Sanh et al. 2019) using the DEBAGREEMENT comment-reply agreement and disagreement data set (Pougué-Biyong et al. 2021), a set of 43 thousand comment-reply pairs from political and climate subreddits, each labeled as either disagreement, neutral, or agreement. Because we are only concerned with the presence of disagreement, constructing a binary classifier based on the three class probabilities results in $f_{1}$ scores of 0.806 and 0.716 for disagreement and lack of disagreement (i.e., neutral or agreement) respectively. Performance details of our fine-tuned DistilBERT model are in Table S4.1.
@@
4.2.4 Obstructionist Claims about Climate Change
To determine whether tweets contain obstructionist claims about climate change, we used the Augmented CARDS (Computer Assisted Recognition of Denial & Skepticism) model presented in (Rojas et al. 2024). The Augmented CARDS contains an initial binary classifier layer based on a DeBERTa neural language model (He et al. 2020), designed to detect tweets as contrarian or otherwise. What constitutes contrarian in the Augmented CARDS model falls under five main categories: 1) global warming is not happening, 2) humans are not causing global warming, 3) climate change is not bad, 4) climate solutions will not work, and 5) climate movement or science are unreliable.
The Augmented CARDS model was trained on an English-language climate change Twitter data set, in addition to the the data set comprising paragraphs extracted from sources known for their contrarian climate content on which the original CARDS model was trained (Coan et al. 2021).
Statistical Analysis
4.3.1 Linear Probability Model and Fractional Logistic Regression Model
Model Sample Size
Linear Probability Model (Toxicity)
Climate General 14,640,829
Climate General FE 14,640,829
Climate Science 109,447
Reddit 756,773
International Organizations 37,010,541
Linear Probability Model (Disagreement)
Climate General 753,812
Climate Science 71,789
Cox Proportional Hazards Model*
Climate General 142,876
Climate Science 22,775
Table 5: Sample sizes for all models. The reported sample size for the Cox proportional hazards models are numbers of independent conversation chains, i.e., the number of samples used in each Cox bootstrap iteration, without consideration of how many posts each conversation chain contains.
We fitted all models in R (R Core Team 2024) with the fixest package (Bergé 2018). Confidence intervals were obtained using the marginaleffects package (Arel-Bundock, Greifer and Heiss 2024). All model sample sizes are reported in Table 5.
@@ 4.3.1 Linear Probability Model and Fractional Logistic Regression Model
We fitted separate linear probability models of climate post toxicity for all three of our data sets – general climate tweets, climate science tweets, and climate Reddit posts. We focused on the COVID-19 period after the initial incivility escalation (August 14, 2020–August 26, 2021) as detected by the BEAST models. In all three models, we included day fixed effects. In all models, we clustered the standard errors at the day level, and also at the subreddit level for the Reddit analysis. We additionally included account fixed effects and account clustering for the general climate tweets data set in a separate model. Fitting all models using fractional logistic regression (Papke and Wooldridge 2008) yielded substantively similar results, which we show in Fig. S6.1.
@@
We fitted separate linear probability models of response disagreement for both of our climate Twitter data sets during the post-escalation COVID-19 period. In both models, we included day fixed effects, and clustered the standard errors at the day, conversation, and initial tweet levels. Results from fractional logistic regression models are substantively similar, which we show in Fig. S6.1.
@@
We fitted a model for tweet toxicity using our general climate Twitter data set for our entire observation period (February 1, 2019–August 26, 2021). In the model, we interacted the international organization keyword term with every quarter in the data (Q1 2019–Q3 2021), then compared the coefficient of the international organization term from each quarter to the pre-pandemic baseline (Q4 2019). We computed standard errors both with and without accounting for clustering. To ensure the increased incivility during the pandemic is not directly due to COVID-19 tweets, we removed them from this analysis, but our results are robust to whether these tweets are included, which we show in Fig. S8.1.
4.3.2 Cox Proportional Hazards Model
To understand the effect of mentions of COVID-19 and other related topics on the onset of toxicity in the course of climate conversations, we constructed an ensemble of Cox proportional hazard models in R (R Core Team 2024) using the survival package (Therneau 2024). Specifically, we modeled the first occurrence of a toxic reply in a Twitter thread, or consecutive chain of tweets and replies, from the post-escalation COVID-19 period. The analysis is based on a resampling method. In each realization, one thread from each conversation tree is selected at random to ensure tweets closer to the root of the conversation are not over-represented in the analysis, as they appear in more threads than tweets closer to the leaves. All model sample sizes are reported in Table 5.
@@ 4.3.2 Cox Proportional Hazards Model
Our models performed adequately in terms of meeting the Schoenfeld test for the proportional hazards assumption (Therneau and Grambsch 2000). Full details of how we implemented the Cox proportional hazards model, including results from the Schoenfeld test, are in supplementary text. We show that our results are robust to alternative specifications of what constitutes an English conversation in Fig. S7.1.
@@
We used Twitter’s Academic API to collect our Twitter data. To collect the general climate tweets, we queried the API for tweets with hashtags and keywords from Table S1.1 for the February 1, 2019–August 26, 2021 period (missed data from February 4, 2021). We excluded tweets containing “political climate” and “Percip:”. The latter is frequently used by bots that automatically tweet regional weather forecast. This yielded 38.4 million English tweets.
To identify all tweets that reference (through hyperlinks) climate science publications, we first used Altmetric’s Explorer API to filter scientific publications by whether their titles contained one of our keywords (Table S1.1), then used Altmetric’s Details Page API to obtain IDs of all tweets that referenced these publications. We used Twitter’s Academic API to collect these tweets. This yielded 258,000 English tweets.
@@
All Twitter data sets were subset to only tweets tagged as English by Twitter and flagged as containing English by Perspective’s multilingual detector. Table S1.2 shows the distribution of the number of users that were active before and after the COVID-19 onset.
@@
Only during the pandemic 2.1M 37K
@@
For our Reddit analysis, we used the Reddit Climate Change Dataset provided by SocialGrep (2022). This data set consists of approximately 4.6 million Reddit comments posted before September 2022 that contain “climate change”. We removed comments posted outside our time window and any comments where Perspective API did not listed English among the languages present. We also removed comments containing phrases indicating that they were written by bots. Specifically, we filtered the comments on a set of common phrases used by bots on Reddit to perform automatic tasks, e.g., “I am a bot” or “This action was performed automatically” (Table S2.1), as well as excluding all comments from the subredditsummarybot because it mostly consist of automatic submissions. Our final Reddit data set contains 2.1 million comments.
@@
Table S2.1: Phrases used for filtering automatic Reddit comments.
@@
Table S3.1: Keywords used for determining mentions of different topics in tweet texts.
@@
We analyzed the general climate conversations and climate science conversations separately. For both sets of analysis, we used all English conversations August 14, 2020 onward from the respective data sets.
@@
  • Tweets with toxicity probability that are not able to be estimated by Perspective API and all subsequent replies in its thread are removed, because we are unable to assess how they affect subsequent replies. (The Perspective API only supports the following languages: Arabic, Chinese, Czech, Dutch, English, French, German, Hindi (including when written with Latin alphabet), Indonesian, Italian, Japanese, Korean, Polish, Portuguese, Russian, Spanish, and Swedish; and is unable to parse tweets containing only non-textual material, e.g., images, videos, or links.) The Perspective API returned NA values for 3.4% of general climate tweets and 6.5% of climate science tweets.
@@
  • If no toxic tweet was found in the entire chain, the thread was treated as right-censored.
  • If a tweet was deleted or made private somewhere along the chain, the thread was likewise treated as right-censored. In terms of systematic deletions that could impact our data, while it is difficult to know the impact of Twitter’s moderation policies, it was reported that as of March 2021 (i.e., approximately seven months before our data collection started), Twitter had deleted only approximately 8400 tweets for COVID-19 misinformation (Twitter 2021).
@@
As threads from the same conversation are not independent from one another due to sharing tweets higher up in the conversation (i.e., closer to the root tweet), we based our analysis on a resampling method. In each realization, we randomly sampled one thread from each conversation. This means that no two threads in the same realization come from the same conversation, and therefore do not shared any tweets. This resampling eliminates the dependence between chains from the same conversation, and also ensures that the proportional hazard assumption is not violated due to presence of extremely large conversations.
@@
Fauci keywords 1.00 1.00
International root 0.98 0.85
International keywords 1.00 1.00
Table S5.1: Cox proportional hazards assumption assessment. Proportion of bootstraps not failing the Schoenfeld test with $p>0.05$.
@@
Linear probability models often provide an adequate approximation of non-linear models, especially when the range of outcome probabilities are not extreme. Still, we fit additional fractional logistic regression models (Papke and Wooldridge 2008) of toxicity probability and disagreement probability with the same model specifications as what we used for our main analysis (Fig. S6.1) to show that our results are robust to accounting for the bounded nature of probabilities.
Figure S6.1: Robustness check of main results using fractional logistic regression models. A. Average marginal effect estimates for toxicity probability, corresponding to results from Fig. 1.B from the main text. B. Average marginal effect estimates for disagreement probability, corresponding to Fig. 2.B from the main text.
@@
Figure S7.1: Robustness check of main Cox proportional hazards model results using full replies instead of only English ones. Figure corresponds to Fig. 1.C from the main text.
@@
When assessing the impact of COVID-19 on the relationship between anti-internationalist sentiment and toxicity in climate discussions, we examined only tweets that do not directly engage with COVID-19 to ensure our results are not driven purely by the toxicity of COVID-19 content. Still, we show that our results are robust to the inclusion of COVID-19 tweets (Fig. S8.1).
Figure S8.1: Robustness check of toxicity probability of international organization tweets over time when including tweets with overlapping COVID-19 content. The lighter confidence intervals are computed from standard errors clustered at the Twitter account level. Figure corresponds to Fig. 3 from the main text.